4x+2y=5
4x+2y=1
are these equations parallel or perpendicular?

Answer by mananth(16946)

4x+2y=5
write the equations in slope intercept form
2y=-4x+5
/2
y=-2x+5/2
4x+2y=1
2y=-4x+1
/2
y=-2x+1/2
compare with y = mx+b
m= slope = -2 for both equations
so they are parallel